OpenStudy (anonymous):

A test consists of multiple choice questions worth 4 points and essay questions worth 8 points. The test has 22 questions that are worth a total of 100 points. Use a model to determine how many multiple choice questions are on the test.

OpenStudy (anonymous):

The answer choices are: 14 16 19 22

OpenStudy (anonymous):

What I got was x = 10 and y = 12. But I guess thats wrong.

OpenStudy (dumbcow):

m+e = 22 4m + 8e = 100 use substitution e=22-m ->4m + 8(22-m) = 100 -4m = -76 m = 19

OpenStudy (anonymous):

simultaneous eqns , all it is

OpenStudy (anonymous):

so this is wrong? m + e = 22 4m + 8e = 100 m = 10, e = 12

OpenStudy (anonymous):

obviously its wrong

OpenStudy (anonymous):

:/

OpenStudy (anonymous):

sub them in, doesnt satisfy second eqn

OpenStudy (anonymous):

ah! I mean. m = 15, e = 5. I don't know where I got 10 and 12 from. ._.

OpenStudy (anonymous):

nvm doesn't work for first equation. JEEZ.

OpenStudy (anonymous):

Thanks for the help. I really do appreciate it. ;D